# Pre-Chirp-Domain Index Modulation for Full-Diversity Affine Frequency Division Multiplexing towards 6G

### v3

@@ I Introduction
The beyond fifth-generation (B5G) and sixth-generation (6G) wireless networks are envisioned to deliver ultra-reliable, high data rate, and low-latency communications for high–speed mobile scenarios, including low-earth-orbit (LEO) satellite, high-mobility railway, unmanned aerial vehicles (UAV) and vehicle-to-vehicle (V2V) communications [2, 3, 4, 5]. These scenarios inevitably suffer from severe Doppler shifts, which can cause time-frequency doubly selective fading (i.e., doubly dispersive channel) by involving the multi-path effects. This makes the existing modulation formats, like the main-stream orthogonal frequency division multiplexing (OFDM) in 4G/5G standards, no longer suitable for next-generation networks [6]. Consequently, it is crucial to develop new waveforms for next-generation communication networks to adapt to the doubly selective channel.
@@
Against this background, affine frequency division multiplexing (AFDM) has been proposed based on the discrete affine Fourier transform (DAFT) [16], which can also combat the time-frequency doubly selective fading, and more importantly has less complexity to implement than OTFS since it requires only one-dimension transformation. DAFT is defined as a generalized discrete form of the discrete Fourier transform (DFT) with a chirp-like basis specified by dual adjustable parameters, termed pre-chirp and post-chirp parameters, respectively. In AFDM, data symbols are multiplexed onto chirp-like subcarriers through DAFT and inverse DAFT (IDAFT), which can separate the doubly dispersive channel into a sparse, quasi-static channel with a comprehensive DD channel representation by appropriately setting the chirp parameters. Therefore, AFDM achieves similar performance to OTFS, and demonstrates superior performance over OFDM and OCDM under doubly selective channels [17].
@@
There has been preliminary literature on AFDM [18, 19, 20]. A low-complexity embedded pilot-aided diagonal reconstruction (EPA-DR) channel estimation scheme was proposed in [18], which calculates the AFDM effective channel matrix directly without estimating the three channel parameters, eliminating the inherently severe inter-Doppler interference. In [19], the authors investigated the AFDM-empowered sparse code multiple access (SCMA) systems to support massive connectivity in high-mobility environments. An AFDM-based integrated sensing and communications (ISAC) system was studied in [20], demonstrating that the AFDM-ISAC system can maintain excellent sensing performance even under significant Doppler shifts. The existing literature mostly explored the channel estimation, multiple access and ISAC issues under the classical AFDM architecture, whilst studies regarding further optimization/enhancement of the AFDM waveform are still at their infancies.
@@
One promising research direction is to incorporate the index modulation (IM) philosophy for spectral and energy efficiency improvement [21, 22], which conveys energy-free bits through the activation patterns of transmit entities, e.g., subcarriers [23], time slots [24], pulse positions [25], antennas [26], etc. In [27, 28], Tao et al. presented an IM-assisted scheme, which conveys energy-free information bits through the activation patterns of the subsymbols in the DAFT-domain, verifying that index bits have stronger diversity protection than modulation bits. A multicarrier system using the activation patterns of AFDM chirp subcarriers as indices was developed in [29], which indicates the potential of IM-assisted AFDM technology in enhancing bit error rate (BER) and energy efficiency performance. However, existing research has concentrated on post-chirp parameters in AFDM, with less attention paid to the considerable flexibility and degrees of freedom (DoFs) that the pre-chirp parameter offers.
@@
The rest of the paper is organized as follows. In Section II, the AFDM system model is introduced. Section III details the proposed AFDM-PIM scheme. The performance analysis of AFDM-PIM under doubly dispersive channels is presented in Section IV, which includes the PEP and diversity analysis. The pre-chirp parameter optimization is provided in Section V. The simulation results and discussions are offered in Section VI, and Section VII draws the conclusions.
@@
Figure 1: The block diagram of AFDM system.
@@ II AFDM System Model
The general system model of AFDM is presented in Fig. 1. For clarity, we provide a concise review of the fundamental concept of AFDM [17]. The transmitted bit stream is initially mapped onto a symbol vector, denoted as $\mathbf{x}_{A}\!=\!\left[x_{A}[0],x_{A}[1],\ldots,x_{A}[N-1]\right]^{\mathrm{T}}\!\in\!\mathbb{C}^{N\times 1}$, comprising $N$ $M$-ary phase shift keying (PSK) symbols in the DAFT domain. The resulting signals are then converted to time-domain representations with an $N$-point IDAFT, formulated as
@@
where $\bar{m}$ is the indices in the DAFT domain. In matrix form, the received AFDM signals can be further written as
$\mathbf{y}_{A}=\sum_{p=1}^{P}h_{p}\mathbf{A}\Gamma_{\textnormal{CPP}_{p}}\bm{\Delta}_{\nu_{p}}\bm{\Pi}^{d_{p}}\mathbf{A}^{\mathrm{H}}\mathbf{x}_{A}+\mathbf{w},$ (6)
@@
and $\Gamma_{\text{CPP}_{p}}=\operatorname{diag}\big{(}\omega_{p,0},\omega_{p,1},\ldots,\omega_{p,N}\big{)}$ is the $N\times N$ diagonal matrix for CPP with
@@ III-A Orthogonality Analysis of AFDM Subcarriers
Figure 2: Transceiver structure of the proposed AFDM-PIM scheme.
TABLE I: Mapping rule between the index bits and the PCPs in the Case of $N_{c}=4$ and $\lambda=4$.
@@ III-B Transmitter
Consider the same AFDM symbol vector comprising $N$ $M$-ary constellation symbols in the DAFT domain as in Section II. At the transmitter, unlike classical AFDM, the $N$ AFDM subcarriers are divided into $G$ groups, with each group comprising $N_{c}\!=\!N/G$ chirp subcarriers. The total $B$ information bits are split into $G$ parallel streams of $b\!=\!B/G$ bits for each subcarrier group. Each $b$-bit stream is further segmented into $b_{1}$ symbol bits and $b_{2}$ index bits, i.e., $b\!=\!b_{1}+b_{2}$. Within the $g$-th group ($1\leq g\leq G$), the $b_{1}\!=\!N_{c}\log_{2}(M)$ symbol bits are conveyed by $N_{c}$ $M$-ary symbols, denoted as $\mathbf{x}^{g}\!=\![x^{g}[0],x^{g}[1],\ldots,x^{g}[N_{c}-1]]^{\mathrm{T}}\!\in\!\mathbb{C}^{N_{c}\times 1}$. On the other hand, each subcarrier is assigned with a unique $c_{2}$ value from a finite alphabet of $\lambda$ legitimate $c_{2}$ realizations, i.e., $\mathcal{P}_{c}\!=\!\left\{{c_{2}^{(1)},c_{2}^{(2)},\ldots,c_{2}^{(\lambda)}}\right\}$. Specifically, the pre-chirping pattern (PCP) of the $c_{2}$ values in the $g$-th group, denoted by $\mathrm{P}_{c_{2}}^{g}\!=\![c_{2,N_{c}\left(g-1\right)},c_{2,N_{c}\left(g-1\right)+1},\ldots,c_{2,N_{c}g-1}]^{\mathrm{T}}\!\in\!\mathbb{C}^{N_{c}\times 1}$ is determined by the $b_{2}$ index bits according to the pre-defined relationship between the index-bit stream and the permutations of $N_{c}$ elements from $\mathcal{P}_{c}$, where $c_{2,m}\!\in\!\mathcal{P}_{c}$ represents the pre-chirp parameter of the $m$-th subcarrier. $\mathrm{P}_{c_{2}}\!=\!\left[\mathrm{P}_{c_{2}}^{1},\mathrm{P}_{c_{2}}^{2},\ldots,\mathrm{P}_{c_{2}}^{G}\right]$ represents the PCP for all the groups (PCPG), and all the possible PCPGs are combined into a set, termed $\mathcal{S}_{p}$. Table I exemplifies the mapping rule in the case of $N_{c}=4$ and $\lambda=4$. Every $b_{2}$ index bits correspond to one row in Table I. Hence, aside from the classical PSK modulation, additional $b_{2}$ information bits can be implicitly conveyed by the indices of $\mathrm{P}_{c_{2}}^{g}$, and
@@
$\bm{\Lambda}_{c_{2}}=\operatorname{diag}\left(e^{-\textsf{j}2\pi c_{2,n}n^{2}},n=0,1,\ldots,N-1\right),$ (16)
@@ III-C Receiver
The received time-domain signals after removing the CPP can be written as
@@
$\mathbf{r}=\mathbf{H}\mathbf{s}+\mathbf{w}=\sum_{p=1}^{P}h_{p}\Gamma_{\textnormal{CPP}_{p}}\bm{\Delta}_{\nu_{p}}\mathbf{\Pi}^{d_{p}}\mathbf{s}+\mathbf{w}_{r},$ (19)
where $\mathbf{w}_{r}\!=\![w_{r}[0],w_{r}[1],\ldots,w_{r}[N-1]]^{\mathrm{T}}$ is the time-domain noise vector. As given in Section II, $\bm{\Delta}_{\nu_{p}}$ represents the Doppler effect, $\mathbf{\Pi}$ is the forward cyclic-shift matrix with $\mathbf{\Pi}^{d_{p}}$ modeling the delay extension, and $\Gamma_{\mathrm{CPP}_{p}}$ is the effective CPP matrix.
@@ IV-A Error Performance Analysis
where $\mathbf{\Phi}(\mathbf{x})\!=\![\mathbf{H}_{1}\mathbf{x},\mathbf{H}_{2}\mathbf{x},\ldots,\mathbf{H}_{P}\mathbf{x}]\!\in\!\mathbb{C}^{N\times P}$ and $\mathbf{h}\!=\![h_{1},h_{2},\ldots,h_{P}]^{\rm T}\!\in\!\mathbb{C}^{P\times 1}$. The conditional PEP (CPEP) between $\mathbf{\Phi}(\mathbf{x})$ and its estimate $\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})$ can be calculated as
$\hskip 28.45274pt=\Pr\left(\mathbf{\chi}>\big{\|}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x})\big{)}\mathbf{h}\big{\|}^{2}\right),$ (32)
where $\chi\!=\!\mathbf{w}^{\rm H}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}\mathbf{h}\!+\!\mathbf{h}^{\rm H}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}^{\rm H}\mathbf{w}$. Since $\chi$ follows Gaussian distribution with variance $2N_{0}\big{\|}(\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x}))\mathbf{h}\big{\|}^{2}$, the CPEP can be expressed as

### v4

@@ I Introduction
The beyond fifth-generation (B5G) and sixth-generation (6G) wireless networks are envisioned to deliver ultra-reliable, high data rate, and low-latency communications for high–speed mobile scenarios, including low-earth-orbit (LEO) satellite, high-mobility railway, unmanned aerial vehicles (UAV) and vehicle-to-vehicle (V2V) communications [2, 3, 4, 5, 6, 7, 8]. These scenarios inevitably suffer from severe Doppler shifts, which can cause time-frequency doubly selective fading (i.e., doubly dispersive channel) by involving the multi-path effects[9]. This makes the existing modulation formats, like the main-stream orthogonal frequency division multiplexing (OFDM) in 4G/5G standards, no longer suitable for next-generation networks [10]. Consequently, it is crucial to develop new waveforms for next-generation communication networks to adapt to the doubly selective channel.
@@
Against this background, affine frequency division multiplexing (AFDM) has been proposed based on the discrete affine Fourier transform (DAFT) [20], which can also combat the time-frequency doubly selective fading, and more importantly has less complexity to implement than OTFS since it requires only one-dimension transformation. DAFT is defined as a generalized discrete form of the discrete Fourier transform (DFT) with a chirp-like basis specified by dual adjustable parameters, termed pre-chirp and post-chirp parameters, respectively. In AFDM, data symbols are multiplexed onto chirp-like subcarriers through DAFT and inverse DAFT (IDAFT), which can separate the doubly dispersive channel into a sparse, quasi-static channel with a comprehensive DD channel representation by appropriately setting the chirp parameters. Therefore, AFDM achieves similar performance to OTFS, and demonstrates superior performance over OFDM and OCDM under doubly selective channels [21], making AFDM a promising waveform for future mobile communication systems [22].
@@
There has been preliminary literature on AFDM [23, 24, 25, 26, 27, 28]. A low-complexity embedded pilot-aided diagonal reconstruction (EPA-DR) channel estimation scheme was proposed in [23], which calculates the AFDM effective channel matrix directly without estimating the three channel parameters, eliminating the inherently severe inter-Doppler interference. In [24], the authors investigated the AFDM-empowered sparse code multiple access (SCMA) systems to support massive connectivity in high-mobility environments. An AFDM-based integrated sensing and communications (ISAC) system was studied in [25], demonstrating that the AFDM-ISAC system can maintain excellent sensing performance even under significant Doppler shifts. In [26], the authors proposed a bistatic sensing-aided channel estimation (SACE) scheme, which illustrates the capacity of AFDM to attain optimal sensing resolution and communication performance under reasonable parameter selection. A DFT-based modulation and demodulation technique for AFDM was studied in [27], with the potential to facilitate the adoption of AFDM for future deployments. The authors in [28] showed that utilizing a single pilot symbol achieves nearly the same sensing performance as employing the complete AFDM frame and proposed a low-complexity self-interference cancellation scheme for monostatic radar. The existing literature mostly explored the channel estimation, multiple access and ISAC issues under the classical AFDM architecture, whilst studies regarding further optimization/enhancement of the AFDM waveform are still at their infancy.
One promising research direction is to incorporate the index modulation (IM) philosophy for spectral and energy efficiency improvement [29, 30, 32, 31], which conveys energy-free bits through the activation patterns of transmit entities, e.g., subcarriers [33], time slots [34], pulse positions [35], antennas [36], constellation modes [37], etc. IM-aided systems provide unique advantages, including flexible system architectures, high spectral efficiency (SE), and low hardware complexity, by leveraging additional energy-free information bits compared to conventional modulation systems.
@@
In [38, 39], Tao et al. presented an IM-assisted scheme, which conveys energy-free information bits through the activation patterns of the subsymbols in the DAFT-domain, verifying that index bits have stronger diversity protection than modulation bits. A multicarrier system using the activation patterns of AFDM chirp subcarriers as indices was developed in [40], which indicates the potential of IM-assisted AFDM technology in enhancing bit error rate (BER) and energy efficiency performance. However, existing research has concentrated on post-chirp parameters in AFDM, with little attention paid to the considerable flexibility and degrees of freedom (DoFs) that the pre-chirp parameter offers. Although the methods of index modulation based on the value of the pre-chirp parameter were discussed in [41], it focused more on reducing the complexity of the demodulation algorithm at the receiver, and no complete design method for the value of different pre-chirp parameters was proposed in this study.
@@
The rest of the paper is organized as follows. In Section II, the AFDM system model is introduced. Section III details the proposed AFDM-PIM scheme. The performance analysis of AFDM-PIM under doubly dispersive channels is presented in Section IV, which includes the PEP and diversity analysis. The pre-chirp parameter optimization is provided in Section V. The simulation results and discussions are offered in Section VI, and Section VII draws the conclusions. The notations adopted in this paper are listed in Table I.
TABLE I: Mathematical Notations
Notation Description
@@
$s\sim\mathcal{CN}(0,\sigma^{2})$ Random variable $s$ follows a complex Gaussian distribution with zero mean and variance $\sigma^{2}$.
$x^{*}$ Conjugate of the complex number $x$.
$a\mid b$ $b$ is divisible by $a$ without leaving a remainder.
$\max\left(a,b\right)$ Maximum value of $a$ and $b$.
$\min\left(a,b\right)$ Minimum value of $a$ and $b$.
$\binom{a}{b}$ Number of ways to choose $b$ elements from a set of $a$ elements.
$(\cdot)_{N}$ Modulo $N$ operation.
$\mathbf{X}^{\mathrm{T}}$, $\mathbf{X}^{\mathrm{H}}$ Transpose and Hermitian operations of $\mathbf{X}$, respectively.
$\left\|\mathbf{X}\right\|_{\mathrm{F}}$ Frobenius norm of $\mathbf{X}$.
$\mathbf{I}_{N}$ $N\times N$ identity matrix.
$Q(\cdot)$ Tail distribution function of the standard Gaussian distribution.
$\mathrm{E}(\cdot)$ Expectation operator.
$\mathbb{I}$ Set of irrational numbers.
$\Re(x)$ Real part of the complex number $x$.
$\textsf{j}=\sqrt{-1}$ Imaginary unit (imaginary axis).
$\prod$ Product operator, representing the multiplication of a sequence of terms.
$\textnormal{Tr}[\cdot]$ Trace operator.
@@ II AFDM System Model
The general system model of AFDM is presented in Fig. 1. For clarity, we provide a concise review of the fundamental concept of AFDM [21]. The transmitted bit stream is initially mapped onto a symbol vector, denoted as $\mathbf{x}_{A}\!=\!\left[x_{A}[0],x_{A}[1],\ldots,x_{A}[N-1]\right]^{\mathrm{T}}\!\in\!\mathbb{C}^{N\times 1}$, comprising $N$ $M$-ary amplitude/phase modulations (APM) symbols in the DAFT domain. The resulting signals are then converted to time-domain representations with an $N$-point IDAFT, formulated as
@@
where $\bar{m}$ are the indices in the DAFT domain. In matrix form, the received AFDM signals can be further written as
$\mathbf{y}_{A}=\sum_{p=1}^{P}h_{p}\mathbf{A}\mathbf{\Gamma}_{\textnormal{CPP}_{p}}\bm{\Delta}_{\nu_{p}}\bm{\Pi}^{d_{p}}\mathbf{A}^{\mathrm{H}}\mathbf{x}_{A}+\mathbf{w},$ (6)
@@
and $\mathbf{\Gamma}_{\text{CPP}_{p}}=\operatorname{diag}\big{(}\omega_{p,0},\omega_{p,1},\ldots,\omega_{p,N}\big{)}$ is the $N\times N$ diagonal matrix for CPP with
@@
Figure 2: Transceiver structure of the proposed AFDM-PIM scheme.
@@ III-A Orthogonality Analysis of AFDM Subcarriers
TABLE II: Mapping rule between the index bits and the PCPs in the Case of $N_{c}=4$ and $\lambda=4$.
@@ III-B Transmitter
Consider the same AFDM symbol vector comprising $N$ $M$-ary constellation symbols in the DAFT domain as in Section II. At the transmitter, unlike classical AFDM, the $N$ AFDM subcarriers are divided into $G$ groups, with each group comprising $N_{c}\!=\!N/G$ chirp subcarriers. The total $B$ information bits are split into $G$ parallel streams of $b\!=\!B/G$ bits for each subcarrier group. Each $b$-bit stream is further segmented into $b_{1}$ symbol bits and $b_{2}$ index bits, i.e., $b\!=\!b_{1}+b_{2}$. Within the $g$-th group ($1\leq g\leq G$), the $b_{1}\!=\!N_{c}\log_{2}(M)$ symbol bits are conveyed by $N_{c}$ $M$-ary symbols, denoted as $\mathbf{x}^{g}\!=\![x^{g}[0],x^{g}[1],\ldots,x^{g}[N_{c}-1]]^{\mathrm{T}}\!\in\!\mathbb{C}^{N_{c}\times 1}$. On the other hand, each subcarrier is assigned with a unique $c_{2}$ value from a finite alphabet of $\lambda$ legitimate $c_{2}$ realizations (corresponding to the ‘$c_{2}$ Bank’ in Fig. 2), i.e., $\mathcal{P}_{c}\!=\!\left\{{c_{2}^{(1)},c_{2}^{(2)},\ldots,c_{2}^{(\lambda)}}\right\}$. Specifically, the pre-chirping pattern (PCP) of the $c_{2}$ values in the $g$-th group, denoted by $\mathrm{P}_{c_{2}}^{g}\!=\![c_{2,N_{c}\left(g-1\right)},c_{2,N_{c}\left(g-1\right)+1},\ldots,c_{2,N_{c}g-1}]^{\mathrm{T}}\!\in\!\mathbb{C}^{N_{c}\times 1}$ is determined by the $b_{2}$ index bits according to the pre-defined relationship between the index-bit stream and the permutations of $N_{c}$ elements from $\mathcal{P}_{c}$, where $c_{2,m}\!\in\!\mathcal{P}_{c}$ represents the pre-chirp parameter of the $m$-th subcarrier. $\mathrm{P}_{c_{2}}\!=\!\left[\mathrm{P}_{c_{2}}^{1},\mathrm{P}_{c_{2}}^{2},\ldots,\mathrm{P}_{c_{2}}^{G}\right]$ represents the PCP for all the groups (PCPG), and all the possible PCPGs are combined into a set, termed $\mathcal{S}_{p}$. Table II exemplifies the mapping rule in the case of $N_{c}=4$ and $\lambda=4$. Every $b_{2}$ index bits correspond to one row in Table II. Hence, aside from the APM bits, additional $b_{2}$ information bits can be implicitly conveyed by the indices of $\mathrm{P}_{c_{2}}^{g}$, and
@@
For ease of design and subsequent optimization, we typically select the case of $\lambda=N_{c}$, which implies $b_{2}=\left\lfloor\log_{2}(N_{c}!)\right\rfloor$.
@@
$\bm{\Lambda}_{c_{2}}=\operatorname{diag}\left(e^{-\textsf{j}2\pi c_{2,m}m^{2}},m=0,1,\ldots,N-1\right),$ (16)
@@ III-C Receiver
At the receiver, considering the doubly dispersive channel in (3), the received time-domain signals after removing the CPP can be written as
@@
$\mathbf{r}=\mathbf{H}\mathbf{s}+\mathbf{w}=\sum_{p=1}^{P}h_{p}\mathbf{\Gamma}_{\textnormal{CPP}_{p}}\bm{\Delta}_{\nu_{p}}\mathbf{\Pi}^{d_{p}}\mathbf{s}+\mathbf{w}_{r},$ (19)
where $\mathbf{w}_{r}\!=\![w_{r}[0],w_{r}[1],\ldots,w_{r}[N-1]]^{\mathrm{T}}$ is the time-domain noise vector. As given in Section II, $\bm{\Delta}_{\nu_{p}}$ represents the Doppler effect, $\mathbf{\Pi}$ is the forward cyclic-shift matrix with $\mathbf{\Pi}^{d_{p}}$ modeling the delay extension, and $\mathbf{\Gamma}_{\mathrm{CPP}_{p}}$ is the effective CPP matrix.
@@
$\hskip 28.45274pt=\mathrm{Pr}\left(\textnormal{Tr}\left[(\mathbf{y}-\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\mathbf{h})^{\rm H}(\mathbf{y}-\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\mathbf{h})-(\mathbf{y}-\mathbf{\Phi}(\mathbf{x})\mathbf{h})^{\rm H}(\mathbf{y}-\mathbf{\Phi}(\mathbf{x})\mathbf{h})\right]<0\right)$
$\hskip 28.45274pt=\mathrm{Pr}\left(\textnormal{Tr}\left[\mathbf{h}^{\rm H}\left(\mathbf{\Phi}(\mathbf{x})-\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})^{\rm H}\right)\left(\mathbf{\Phi}(\mathbf{x})-\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})^{\rm H}\right)\mathbf{h}\right]-\mathrm{C}<0\right)=\Pr\left(\mathbf{\chi}>\big{\|}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x})\big{)}\mathbf{h}\big{\|}^{2}\right),$ (32)
@@ IV-A Error Performance Analysis
where $\mathbf{\Phi}(\mathbf{x})\!=\![\mathbf{H}_{1}\mathbf{x},\mathbf{H}_{2}\mathbf{x},\ldots,\mathbf{H}_{P}\mathbf{x}]\!\in\!\mathbb{C}^{N\times P}$ and $\mathbf{h}\!=\![h_{1},h_{2},\ldots,h_{P}]^{\rm T}\!\in\!\mathbb{C}^{P\times 1}$. The conditional PEP (CPEP) between $\mathbf{\Phi}(\mathbf{x})$ and its estimate $\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})$ can be calculated as (III-C), shown at the bottom of this page, where $\chi\!=\!\mathbf{w}^{\rm H}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}\mathbf{h}\!+\!\mathbf{h}^{\rm H}\big{(}\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})\!-\!\mathbf{\Phi}(\mathbf{x})\big{)}^{\rm H}\mathbf{w}$. Since $\chi$ follows Gaussian distribution with variance $2N_{0}\big{\|}(\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})-\mathbf{\Phi}(\mathbf{x}))\mathbf{h}\big{\|}^{2}$, the CPEP can be expressed as